### v2

@@ B.1 Constrained optimal rule without functional form restriction
$\max_{g:\mathcal{X}\rightarrow\{0,1\}}E[\Gamma_{i}g(X_{i})]\ s.t.\ E[R_{i}g(X_{i})]<k$
@@
$\max_{g:\mathcal{X}\rightarrow\{0,1\}}E[\gamma(X_{i})g(X_{i})]\ s.t.\ E[r(X_{i})g(X_{i})]$ $<k$
where $\gamma(X_{i})=E[\Gamma_{i}\mid X_{i}]$ and $r(X_{i})=E[R_{i}\mid X_{i}]$.
@@
[…]
@@
By definition of $X^{\ast}$, we have $\frac{\gamma(x)}{r(x)}>c$ for $x\in X^{\ast}\backslash X$ and $\frac{\gamma(x)}{r(x)}<c$ for $x\in X\backslash X^{\ast}$. Also note that $\mu$ is a positive measure. Then the above difference is lower bounded by
@@ B.2 Primitive assumptions for contiguity
[…]
@@
For all policies $g$, $B(g;P_{\theta})$ is twice continuously differentiable in $\theta$ at 00, and the derivatives are bounded from above and away from zero within an open neighborhood $\mathcal{N}_{\theta}$ of zero uniformly over $g\in\mathcal{G}$.
@@
$\log\prod_{i=1}^{n}\frac{p_{h_{n}}}{p_{0}}(A_{i})$ $=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}h^{\prime}\dot{\ell}_{0}(A_{i})-\frac{1}{2}h^{\prime}I_{0}h+o_{P_{0}}(1)$ (25)
@@
where ˙$\dot{\ell}_{0}$ is the score and $I_{P_{0}}=E_{P_{0}}[\dot{\ell}_{0}(A_{i})\dot{\ell}_{0}(A_{i})^{\prime}]$ exists. The convergence in distribution of the log likelihood ratio to a normal with mean equal to $-\frac{1}{2}$ of its variance in (26) implies mutual contiguity $P_{0}^{n}\vartriangleleft\vartriangleright P_{h_{n}}^{n}$ by Le Cam’s first lemma (see Example 6.5 of Vaart (1998)). This proves the first part of the lemma.
@@
for every $n$. This proves the second part of the lemma for $C=\inf_{g\in\mathcal{G}_{0}}\left|h^{\prime}\frac{\partial B(g;P_{0})}{\partial\theta}\right|$ ∎
@@ B.3 Primitive assumptions and proofs for estimation quality
$\widehat{W}_{n}(g)\coloneqq\frac{1}{n}\sum_{i}\Gamma_{i}^{\ast}\cdot g(X_{i}),\ \widehat{B}_{n}(g)\coloneqq\frac{1}{n}\sum_{i}R_{i}^{\ast}\cdot g(X_{i})$
@@
$\Gamma_{i}^{\ast}$ $=\widehat{\gamma}^{Y}(V_{i},1)-\widehat{\gamma}^{Y}(V_{i},0)+\widehat{\alpha}(V_{i},D_{i})\cdot\left(Y_{i}-\widehat{\gamma}^{Y}(V_{i},D_{i})\right)$
$R_{i}^{\ast}$ $=\widehat{\gamma}^{Z}(V_{i},1)+\frac{D_{i}}{\widehat{p}(V_{i})}\cdot\left(Z_{i}-\widehat{\gamma}^{Z}(V_{i},D_{i})\right)$
@@
Recall that $\widehat{W}_{n}(g)$ and $\widehat{B}_{n}(g)$ are supposed to approximate net benefit $\Gamma$ and net excess cost $R$ of Medicaid eligibility. I provide more precise definitions for $\Gamma$ and $R$ as the primitive assumptions are stated in terms of their components.
@@
Let $C(1)$ be the (potential) cost to the government when one is given Medicaid eligibility, and $C(0)$ be the (potential) cost to the government when one is not given Medicaid eligibility. Note that $C(0)=0$ by construction. Even when given eligibility, one might not enroll and thus incur zero cost to the government. So given the current expenditure is \$6,000 per enrollee, the implied expenditure per capita under eligibility policy $g(X)$ is actually \$$\$6,000\cdot E\left[M(1)g(X)\right]$. The reason is that the expected enrollment rate is only $Pr\{M(1)=1\mid X\in g\}$. So the per capita excess cost of Medicaid eligibility policy $g(X)$ relative to the current level is
$R=C(1)-\$6,000\cdot M(1).$
We only observe the actual cost ($C_{i}=D_{i}C_{i}(1)+(1-D_{i})C_{i}(0)$) and the actual Medicaid enrollment ($M_{i}$) in OHIE. We calculate $Z_{i}=C_{i}-\$6,000\cdot M_{i}$.
@@
Independent characteristics: $Pr\{D_{i}=1\mid V_{i},X_{i}\}=Pr\{D_{i}=1\mid V_{i}\}$
Unconfoundedness: $(Y(1),Y(0),C(1),M(1))\perp D_{i}|V_{i}$.
@@ B.3.1 Uniform convergence of $\widehat{W}_{n}(\cdot)$ and $\widehat{B}_{n}(\cdot)$
$\sqrt{n}\left(\frac{1}{n}\sum_{i}\Gamma_{i}^{\ast}\cdot g(X_{i})-E_{P}[\Gamma\cdot g(X_{i})]\right)$
@@
$\frac{1}{\sqrt{n}}\sum_{i}(\Gamma_{i}^{\ast}-\widetilde{\Gamma}_{i})\cdot g(X_{i})+\sqrt{n}\left(\frac{1}{n}\sum_{i}\widetilde{\Gamma}_{i}\cdot g(X_{i})-E_{P}[\Gamma\cdot g(X_{i})]\right).$ (27)
Here $\widetilde{\Gamma}_{i}$ are the theoretical analogs
$\widetilde{\Gamma}_{i}$ $=\gamma^{Y}(V_{i},1)-\gamma^{Y}(V_{i},0)+\alpha(V_{i},D_{i})\cdot\left(Y_{i}-\gamma^{Y}(V_{i},D_{i})\right)$
@@
$\widetilde{R}_{i}=\gamma^{Z}(V_{i},1)+\frac{D_{i}}{p(V_{i})}\cdot\left(Z_{i}-\gamma^{Z}(V_{i},D_{i})\right).$
@@
$\widehat{\Sigma}^{B}(g,g^{\prime})$ $=\frac{1}{n}\sum_{i}(\widetilde{R}_{i})^{2}\cdot g(X_{i})\cdot g^{\prime}(X_{i})-\left(\frac{1}{n}\sum_{i}\widetilde{R}_{i}\cdot g(X_{i})\right)\cdot\left(\frac{1}{n}\sum_{i}\widetilde{R}_{i}\cdot g^{\prime}(X_{i})\right)$
@@
$\mathcal{F}^{W}$ =$=\{\widetilde{\Gamma}_{i}\cdot g(X_{i}):g\in\mathcal{G}\}$
$\mathcal{F}^{B}$ =$=\{\widetilde{R}_{i}\cdot g(X_{i}):g\in\mathcal{G}\}$
@@
$E_{P}[\widetilde{\Gamma}_{i}\cdot g(X_{i})]$ $=W(g;P)$
$E_{P}[\widetilde{R}_{i}\cdot g(X_{i})]$ $=B(g;P)$
@@
$\sqrt{n}\cdot(\frac{1}{n}\sum_{i}\widetilde{\Gamma}_{i}\cdot g(X_{i})-W(g;P))$
@@
$\sqrt{n}\cdot(\frac{1}{n}\sum_{i}\widetilde{R}_{i}\cdot g(X_{i})-B(g;P))$
@@
$\sup_{P\in\mathcal{P}}\sup_{g\in\mathcal{G}}\left|\frac{1}{\sqrt{n}}\sum_{i}(\Gamma_{i}^{\ast}-\widetilde{\Gamma}_{i})\cdot g(X_{i})\right|$ $\rightarrow_{p}0$
$\sup_{P\in\mathcal{P}}\sup_{g\in\mathcal{G}}\left|\frac{1}{\sqrt{n}}\sum_{i}(R_{i}^{\ast}-\widetilde{R}_{i})\cdot g(X_{i})\right|$ $\rightarrow_{p}0$
@@ B.3.2 Proofs of auxiliary lemmas
Proof of Lemma B.3. Under Assumption B.4, we prove each $\widetilde{\Gamma}_{i}$ is an conditionally unbiased estimate for $\Gamma$:
$E_{P}[\widetilde{\Gamma}_{i}g(X_{i})]=E_{P}[E_{P}[\widetilde{\Gamma}_{i}\mid X_{i}]g(X_{i})]=E_{P}[E_{P}[\Gamma\mid X_{i}]g(X_{i})]=W(g;P).$
We focus on $E_{P}[\widetilde{\Gamma}_{i}\mid X_{i}]=E_{P}[E_{P}[\widetilde{\Gamma}_{i}\mid V_{i},X_{i}]\mid X_{i}]$. However, conditional on $V_{i}$, by unconfoundedness and strict overlap we have
$E_{P}[\widetilde{\Gamma}_{i}\mid V_{i},X_{i}]$ $=E[Y_{i}\mid V_{i},X_{i},D_{i}=1]-E[Y_{i}\mid V_{i},X_{i},D_{i}=0]$
$=E[Y_{i}(1)-Y_{i}(0)\mid V_{i},X_{i}]=E[\Gamma_{i}\mid V_{i},X_{i}]$
@@
$E[\widetilde{\Gamma}_{i}\mid V_{i},X_{i}]$ $=E\left[\gamma^{Y}(V_{i},1)-\gamma^{Y}(V_{i},0)\mid V_{i},X_{i}\right]+E\left[\alpha(V_{i},D_{i})\cdot\left(Y_{i}-\gamma^{Y}(V_{i},D_{i})\right)\mid V_{i},X_{i}\right]$
$=E\left[Y_{i}\mid V_{i},D_{i}=1\right]-E\left[Y_{i}\mid V_{i},D_{i}=0\right]+$
$E\left[Y_{i}(1)-Y_{i}(0)\mid V_{i},X_{i}\right]-(E\left[Y_{i}\mid V_{i},D_{i}=1\right]-E\left[Y_{i}\mid V_{i},D_{i}=0\right])$
$=E[\Gamma\mid V_{i},X_{i}]$
Specifically, we expand $E[\alpha(V_{i},D_{i})Y_{i}\mid V_{i},X_{i}]$
$=E\left[\frac{Y_{i}}{p(V_{i})}\mid D_{i}=1,V_{i},X_{i}\right]\cdot Pr\{D_{i}=1\mid V_{i},X_{i}\}-E\left[\frac{Y_{i}}{1-p(V_{i})}\mid D_{i}=0,X_{i}\right]\cdot Pr\{D_{i}=0\mid V_{i},X_{i}\}$
$=E\left[D_{i}Y_{i}(1)+(1-D_{i})Y_{i}(0)\mid D_{i}=1,V_{i},X_{i}\right]-E\left[D_{i}Y_{i}(1)+(1-D_{i})Y_{i}(0)\mid D_{i}=0,V_{i},X_{i}\right]$
$=E\left[Y_{i}(1)\mid V_{i},X_{i}\right]-E\left[Y_{i}(0)\mid V_{i},X_{i}\right]$
@@
$E\left[\alpha(V_{i},D_{i})\cdot\gamma^{Y}(V_{i},D_{i})\mid V_{i},X_{i}\right]=E\left[Y_{i}\mid V_{i},D_{i}=1\right]-E\left[Y_{i}\mid V_{i},D_{i}=0\right]$
Similar argument holds for $E_{P}[\widetilde{R}_{i}\cdot g(X_{i})]$, with the only modification:
$E[\frac{D_{i}}{p(V_{i})}C_{i}\mid V_{i},X_{i}]=E\left[\frac{C_{i}}{p(V_{i})}\mid D_{i}=1,V_{i},X_{i}\right]\cdot Pr\{D_{i}=1\mid V_{i},X_{i}\}$
$=E[C_{i}(1)\mid V_{i},X_{i}]$
@@
$\mathcal{F}^{U}$ =$=\{U_{i}\cdot g(X_{i}):g\in\mathcal{G}\}$
@@
Proof of Lemma B.6. We focus on the deviation in $\Gamma_{i}^{\ast}-\widetilde{\Gamma}_{i}$. The deviation in $R_{i}^{\ast}-\widetilde{R}_{i}$ can be proven to vanish in a similar manner. Denote $\Delta\gamma^{Y}(V_{i})=\gamma^{Y}(V_{i},1)-\gamma^{Y}(V_{i},0)$. For any fixed policy $g$, we expand the deviation $\frac{1}{\sqrt{n}}\sum_{i}(\Gamma_{i}^{\ast}-\widetilde{\Gamma}_{i})g(X_{i})$ into three terms
@@
where $\widehat{M}=\frac{1}{n}\sum_{i}b(V_{i},D_{i})b(V_{i},D_{i})^{\prime}.$ It converges in probability to a fixed matrix $M=E[b(V_{i},D_{i})b(V_{i},D_{i})^{\prime}]$. So the in-sample $L_{2}$ error from the estimated CEF vanishes at the rate of $n^{-1/2}$.

### v3

@@ B.1 Constrained optimal rule without functional form restriction
$\max_{g:\mathcal{X}\rightarrow\{0,1\}}\mathbb{E}[\tau_{i}g(X_{i})]\text{ subject to }\mathbb{E}[C_{i}g(X_{i})]<k$
@@
$\max_{g:\mathcal{X}\rightarrow\{0,1\}}\mathbb{E}[\gamma(X_{i})g(X_{i})]\text{ subject to }\mathbb{E}[r(X_{i})g(X_{i})]$ $<k$
where $\gamma(X_{i})=\mathbb{E}[\tau_{i}\mid X_{i}]$ and $r(X_{i})=\mathbb{E}[C_{i}\mid X_{i}]$.
@@
$\max_{g:\mathcal{X}\rightarrow\{0,1\}}\int\frac{\gamma(x)}{r(x)}g(x)d\mu\text{ subject to }\int g(x)d\mu=k$
@@
$\int_{X^{\ast}}\frac{\gamma}{r}d\mu-\int_{X}\frac{\gamma}{r}d\mu$ $=\int\frac{\gamma(x)}{r(x)}\mathbf{1}\{x\in X^{\ast}\}d\mu-\int\frac{\gamma(x)}{r(x)}\mathbf{1}\{x\in X\}d\mu$
@@ B.2 Primitive assumptions for contiguity
Assume the class of DGPs $\{P_{\theta}:\theta\in\Theta\}$ has densities $p_{\theta}$ with respect to some measure $\mu$. Assume $P_{\theta}$ is DQM at $P_{0}$ i.e. $\exists\dot{\ell}_{0}$ s.t. $\int[\sqrt{p_{h}}-\sqrt{p_{0}}-\frac{1}{2}h^{\prime}\dot{\ell}_{0}\sqrt{p_{0}}]^{2}d\mu=o(\left\|h^{2}\right\|)$ for $h\rightarrow 0$.
@@
For all policies $g$, $B(g;P_{\theta})$ is twice continuously differentiable in $\theta$ at 0, and the derivatives are bounded from above and away from zero within an open neighborhood $\mathcal{N}_{\theta}$ of zero uniformly over $g\in\mathcal{G}$.
@@
$\log\prod_{i=1}^{n}\frac{p_{h_{n}}}{p_{0}}(A_{i})$ $=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}h^{\prime}\dot{\ell}_{0}(A_{i})-\frac{1}{2}h^{\prime}I_{0}h+o_{P_{0}}(1)$
@@
where $\dot{\ell}_{0}$ is the score and $I_{P_{0}}=E_{P_{0}}[\dot{\ell}_{0}(A_{i})\dot{\ell}_{0}(A_{i})^{\prime}]$ exists. The convergence in distribution of the log likelihood ratio to a normal with mean equal to $\frac{1}{2}$ of its variance in (16) implies mutual contiguity $P_{0}^{n}\vartriangleleft\vartriangleright P_{h_{n}}^{n}$ by Le Cam’s first lemma (see Example 6.5 of Vaart (1998)). This proves the first part of the lemma.
@@
for every $n$. This proves the second part of the lemma for $c=\inf_{g\in\mathcal{G}_{0}}\left|h^{\prime}\frac{\partial B(g;P_{0})}{\partial\theta}\right|$ . ∎
@@ B.3 Primitive assumptions and proofs for estimation quality
$\widehat{W}_{n}(g)\coloneqq\frac{1}{n}\sum_{i}\tau_{i}^{\ast}\cdot g(X_{i}),\ \widehat{B}_{n}(g)\coloneqq\frac{1}{n}\sum_{i}C_{i}^{\ast}\cdot g(X_{i})$
@@
$\tau_{i}^{\ast}$ $=\widehat{\gamma}^{Y}(V_{i},1)-\widehat{\gamma}^{Y}(V_{i},0)+\widehat{\alpha}(V_{i},D_{i})\cdot\left(Y_{i}-\widehat{\gamma}^{Y}(V_{i},D_{i})\right)$
$C_{i}^{\ast}$ $=\widehat{\gamma}^{Z}(V_{i},1)+\frac{D_{i}}{\widehat{p}(V_{i})}\cdot\left(Z_{i}-\widehat{\gamma}^{Z}(V_{i},D_{i})\right)$
@@
Recall that $\widehat{W}_{n}(g)$ and $\widehat{B}_{n}(g)$ are supposed to approximate net benefit $\tau$ and net excess cost $C$ of Medicaid eligibility. I provide more precise definitions for $\tau$ and $C$ as the primitive assumptions are stated in terms of their components.
@@
Let $H(1)$ be the (potential) health care utilization when one is given Medicaid eligibility, and $H(0)$ be the (potential) health care utilization when one is not given Medicaid eligibility. Even when given eligibility, one might not enroll and thus incur zero cost to the government. So the per capita cost of Medicaid eligibility policy $g(X)$ is
$\mathbb{E}_{P}[C\cdot g(X)]\text{ where }C=M(1)\cdot H(1).$
We only observe the actual health care utilization ($H_{i}=D_{i}H_{i}(1)+(1-D_{i})H_{i}(0)$) and the actual Medicaid enrollment ($M_{i}$) in OHIE. We calculate $Z_{i}=M_{i}\cdot H_{i}$.
@@
Independent characteristics: $\operatorname{Pr}\{D_{i}=1\mid V_{i},X_{i}\}=\operatorname{Pr}\{D_{i}=1\mid V_{i}\}$
Unconfoundedness: $(Y(1),Y(0),H(1),M(1))\perp D_{i}|V_{i}$.
@@ B.3.1 Uniform convergence of $\widehat{W}_{n}(\cdot)$ and $\widehat{B}_{n}(\cdot)$
$\sqrt{n}\left(\frac{1}{n}\sum_{i}\tau_{i}^{\ast}\cdot g(X_{i})-\mathbb{E}_{P}[\tau\cdot g(X_{i})]\right)$
@@
$\frac{1}{\sqrt{n}}\sum_{i}(\tau_{i}^{\ast}-\widetilde{\tau}_{i})\cdot g(X_{i})+\sqrt{n}\left(\frac{1}{n}\sum_{i}\widetilde{\tau}_{i}\cdot g(X_{i})-\mathbb{E}_{P}[\tau\cdot g(X_{i})]\right).$
Here $\widetilde{\tau}_{i}$ are the theoretical analogs
$\widetilde{\tau}_{i}$ $=\gamma^{Y}(V_{i},1)-\gamma^{Y}(V_{i},0)+\alpha(V_{i},D_{i})\cdot\left(Y_{i}-\gamma^{Y}(V_{i},D_{i})\right)$
@@
$\widetilde{C}_{i}=\gamma^{Z}(V_{i},1)+\frac{D_{i}}{p(V_{i})}\cdot\left(Z_{i}-\gamma^{Z}(V_{i},D_{i})\right).$
@@
$\widehat{\Sigma}^{B}(g,g^{\prime})$ $=\frac{1}{n}\sum_{i}(\widetilde{C}_{i})^{2}\cdot g(X_{i})\cdot g^{\prime}(X_{i})-\left(\frac{1}{n}\sum_{i}\widetilde{C}_{i}\cdot g(X_{i})\right)\cdot\left(\frac{1}{n}\sum_{i}\widetilde{C}_{i}\cdot g^{\prime}(X_{i})\right)$
@@
$\mathcal{F}^{W}$ $=\{\widetilde{\tau}_{i}\cdot g(X_{i}):g\in\mathcal{G}\}$
$\mathcal{F}^{B}$ $=\{\widetilde{C}_{i}\cdot g(X_{i}):g\in\mathcal{G}\}$
@@
$\mathbb{E}_{P}[\widetilde{\tau}_{i}\cdot g(X_{i})]$ $=W(g;P)$
$\mathbb{E}_{P}[\widetilde{C}_{i}\cdot g(X_{i})]$ $=B(g;P)$
@@
$\sqrt{n}\cdot(\frac{1}{n}\sum_{i}\widetilde{\tau}_{i}\cdot g(X_{i})-W(g;P))$
@@
$\sqrt{n}\cdot(\frac{1}{n}\sum_{i}\widetilde{C}_{i}\cdot g(X_{i})-B(g;P))$
@@
$\sup_{P\in\mathcal{P}}\sup_{g\in\mathcal{G}}\left|\frac{1}{\sqrt{n}}\sum_{i}(\tau_{i}^{\ast}-\widetilde{\tau}_{i})\cdot g(X_{i})\right|$ $\rightarrow_{p}0$
$\sup_{P\in\mathcal{P}}\sup_{g\in\mathcal{G}}\left|\frac{1}{\sqrt{n}}\sum_{i}(C_{i}^{\ast}-\widetilde{C}_{i})\cdot g(X_{i})\right|$ $\rightarrow_{p}0$
@@ B.3.2 Proofs of auxiliary lemmas
Proof of Lemma B.3. Under Assumption B.4, we prove each $\widetilde{\tau}_{i}$ is an conditionally unbiased estimate for $\tau$:
$\mathbb{E}_{P}[\widetilde{\tau}_{i}g(X_{i})]=\mathbb{E}_{P}[\mathbb{E}_{P}[\widetilde{\tau}_{i}\mid X_{i}]g(X_{i})]=\mathbb{E}_{P}[\mathbb{E}_{P}[\tau\mid X_{i}]g(X_{i})]=W(g;P).$
We focus on $\mathbb{E}_{P}[\widetilde{\tau}_{i}\mid X_{i}]=\mathbb{E}_{P}[\mathbb{E}_{P}[\widetilde{\tau}_{i}\mid V_{i},X_{i}]\mid X_{i}]$. Dropping the subscript $P$ for simplicity, conditional on $V_{i}$, by unconfoundedness and strict overlap we have
$\mathbb{E}[\widetilde{\tau}_{i}\mid V_{i},X_{i}]$ $=\mathbb{E}[Y_{i}\mid V_{i},X_{i},D_{i}=1]-\mathbb{E}[Y_{i}\mid V_{i},X_{i},D_{i}=0]$
$=\mathbb{E}[Y_{i}(1)-Y_{i}(0)\mid V_{i},X_{i}]=\mathbb{E}[\tau_{i}\mid V_{i},X_{i}]$
@@
$\mathbb{E}[\widetilde{\tau}_{i}\mid V_{i},X_{i}]$ $=\mathbb{E}\left[\gamma^{Y}(V_{i},1)-\gamma^{Y}(V_{i},0)\mid V_{i},X_{i}\right]+\mathbb{E}\left[\alpha(V_{i},D_{i})\cdot\left(Y_{i}-\gamma^{Y}(V_{i},D_{i})\right)\mid V_{i},X_{i}\right]$
$=\mathbb{E}\left[Y_{i}\mid V_{i},D_{i}=1\right]-\mathbb{E}\left[Y_{i}\mid V_{i},D_{i}=0\right]+$
$\mathbb{E}\left[Y_{i}(1)-Y_{i}(0)\mid V_{i},X_{i}\right]-(\mathbb{E}\left[Y_{i}\mid V_{i},D_{i}=1\right]-\mathbb{E}\left[Y_{i}\mid V_{i},D_{i}=0\right])$
$=\mathbb{E}[\tau\mid V_{i},X_{i}]$
Specifically, we expand $\mathbb{E}[\alpha(V_{i},D_{i})Y_{i}\mid V_{i},X_{i}]$
$=\mathbb{E}\left[\frac{Y_{i}}{p(V_{i})}\mid D_{i}=1,V_{i},X_{i}\right]\cdot\operatorname{Pr}\{D_{i}=1\mid V_{i},X_{i}\}-\mathbb{E}\left[\frac{Y_{i}}{1-p(V_{i})}\mid D_{i}=0,X_{i}\right]\cdot\operatorname{Pr}\{D_{i}=0\mid V_{i},X_{i}\}$
$=\mathbb{E}\left[D_{i}Y_{i}(1)+(1-D_{i})Y_{i}(0)\mid D_{i}=1,V_{i},X_{i}\right]-\mathbb{E}\left[D_{i}Y_{i}(1)+(1-D_{i})Y_{i}(0)\mid D_{i}=0,V_{i},X_{i}\right]$
$=\mathbb{E}\left[Y_{i}(1)\mid V_{i},X_{i}\right]-\mathbb{E}\left[Y_{i}(0)\mid V_{i},X_{i}\right]$
@@
$\mathbb{E}\left[\alpha(V_{i},D_{i})\cdot\gamma^{Y}(V_{i},D_{i})\mid V_{i},X_{i}\right]=\mathbb{E}\left[Y_{i}\mid V_{i},D_{i}=1\right]-\mathbb{E}\left[Y_{i}\mid V_{i},D_{i}=0\right]$
Similar argument holds for $\mathbb{E}_{P}[\widetilde{C}_{i}\cdot g(X_{i})]$, with the only modification:
$\mathbb{E}[\frac{D_{i}}{p(V_{i})}H_{i}\mid V_{i},X_{i}]=\mathbb{E}\left[\frac{H_{i}}{p(V_{i})}\mid D_{i}=1,V_{i},X_{i}\right]\cdot\operatorname{Pr}\{D_{i}=1\mid V_{i},X_{i}\}$
$=\mathbb{E}[H_{i}(1)\mid V_{i},X_{i}]$
@@
$\mathcal{F}^{U}$ $=\{U_{i}\cdot g(X_{i}):g\in\mathcal{G}\}$
@@
Proof of Lemma B.6. We focus on the deviation in $\tau_{i}^{\ast}-\widetilde{\tau}_{i}$. The deviation in $C_{i}^{\ast}-\widetilde{C}_{i}$ can be proven to vanish in a similar manner. Denote $\Delta\gamma^{Y}(V_{i})=\gamma^{Y}(V_{i},1)-\gamma^{Y}(V_{i},0)$. For any fixed policy $g$, we expand the deviation $\frac{1}{\sqrt{n}}\sum_{i}(\tau_{i}^{\ast}-\widetilde{\tau}_{i})g(X_{i})$ into three terms
@@
where $\widehat{M}=\frac{1}{n}\sum_{i}b(V_{i},D_{i})b(V_{i},D_{i})^{\prime}.$ It converges in probability to a fixed matrix $M=\mathbb{E}[b(V_{i},D_{i})b(V_{i},D_{i})^{\prime}]$. So the in-sample $L_{2}$ error from the estimated CEF vanishes at the rate of $n^{-1/2}$.